### v2

# Nonlinear particle motion and bursty periodic energy deposition in inductively coupled plasmas

Abstract
Two-dimensional, electromagnetic particle-in-cell simulations are employed to study particle kinetics and power deposition in inductively coupled plasmas. We show that in the regime of low frequency ($\sim\mathrm{MHz}$) and low pressure, the motion of electrons can be highly nonlinear in the skin region near the antenna coil. Through most of the RF cycle, the electrons are strongly magnetized, with energy deposition being small. However, for a brief period around the null of the oscillating magnetic field, the electrons demagnetize, causing a jet-like current penetrating into the bulk plasma. During these brief periods the power deposition becomes high, exhibiting a periodic burst nature. Based on kinetic theory, we provide new analytical expressions for the plasma current and energy deposition. A criterion for transition between the newly identified low-frequency, periodic-burst regime and the usual anomalous non-local skin effect regime is proposed and verified using a series of fully kinetic 2D particle-in-cell simulations.
preprint: APS/123-QED
I
Introduction–Energy deposition into electrons by a time-varying electromagnetic field is the primary method of sustaining the plasma in gas discharges for material processing [1, 2], such as Capacitively Coupled Plasma (CCP) [3, 4, 5, 6, 7, 8, 6, 9, 10, 11] and Inductively Coupled Plasma (ICP) discharges [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Driven by industry demands to achieve atomic-scale precision, plasma processing systems tend to operate at low gas pressure (few mTorr) [22, 23, 24, 25, 26, 27, 7, 5, 28, 9, 29, 30, 31], making it crucial to understand the underlying electron dynamics [32, 33]. For low-pressure, high-frequency ($f=13.56~{}\mathrm{MHz}$ or higher) ICP discharges, the skin effect was found to be of the anomalous type [23, 24, 25, 26, 34, 27, 35, 36, 37, 38], meaning that the skin depth, $\delta$, is determined by non-local kinetics as $\delta\sim(v_{th}c^{2}/\omega\omega^{2}_{pe})^{1/3}$ instead of $\delta\sim c/\omega_{pe}$ for the normal skin effect, where $v_{th}$ is the electron thermal velocity, $c$ is the speed of light, $\omega$ is the driving frequency, and $\omega_{pe}$ is the electron plasma frequency. The electron heating in the anomalous regime is mainly due to wave-particle resonance of low-energy electrons, which can be treated similarly to the Landau damping [16, 39, 40, 38, 41, 18, 32, 42]. The anomalous regime had also been called the non-local regime because over one RF period the electrons travel a significant distance $d\sim v_{th}/\sqrt{\omega^{2}+\nu^{2}}>\delta$ [38], where $\nu$ is the collision frequency. As a result, the plasma current is a non-local function of the RF electric field, and the RF field penetrates deeper into the plasma and deposits its energy over a larger depth. Multiple modeling studies have been performed for the ICP discharges, utilizing the non-local kinetic approach [16, 43, 32, 44, 45, 46, 47, 48], to predict the spatial distribution of the plasma current as well as the resulting power deposition.
Although historically ICP discharge systems operated mostly at high frequency [49, 50], it has been recently shown that operating at lower frequency ($\sim 1$MHz) presents certain advantages [51, 52, 53, 54]. At low frequency, the capacitive coupling with the plasma can be reduced [55], the external circuit losses decrease [56, 55], and higher power transfer efficiency can be achieved [52], thereby improving the overall performance of plasma processing systems. A more complicated ICP operation under pulsed waveforms [57, 58, 59, 60, 61, 62, 63] also requires a fundamental understanding of ICP discharge at low frequency. Low-frequency ICP discharges are also utilized as sources of negative ions for generating fast atomic beams in fusion applications, due to improved plasma uniformity [64, 65, 66, 67].
At the same time, due to the lack of a self-consistent kinetic electromagnetic 2D modeling, the physics of low-frequency ICP discharges is not sufficiently understood, because in experiments it is difficult to perform detailed measurements of plasma properties and electromagnetic fields within the skin layer and correspondingly deduce complex particle kinetics [53, 34]. Although it was previously shown that at low frequencies the particle dynamics could become highly nonlinear due to trapping in the RF magnetic field [68, 69], those studies were not based on fully self-consistent kinetic simulations and had to use simplified test-particle analysis. Ref. [70] classified ICP discharges according to their estimated skin depth, assuming that the electron gyro-radius is much larger than the skin depth. However, as will be shown below it is not the case at high coil current where RF magnetic field strongly affects particle motion. Moreover, existing kinetic simulation studies have been limited to one-dimensional [71, 72, 73, 44] or high-frequency ICP discharges [74, 75]. More importantly, to the best of our knowledge, a theory that could explain the particle dynamics and predict the analytical expression for the electron current in a low-frequency inductive discharge has been lacking. Therefore, a comprehensive numerical and analytical study is necessary to understand the relevant particle dynamics and the resulting power deposition in low-frequency ICP discharges.
We demonstrate that a strongly nonlinear skin effect regime with periodic bursty energy deposition is attained at a sufficiently low RF frequency. In this regime, electrons are magnetized and power deposition is low during most of the RF period. Demagnetization occurs when the RF magnetic field passes through zero, and electrons quickly escape area near the antennas to form a current jet propagating into the plasma, causing periodic bursts in the energy deposition. A new kinetic theory is developed and predicts well the plasma current in this bursty regime. To the best of our knowledge, this work presents the first comprehensive 2D PIC simulations of a low frequency ICP discharge.
Method – We deploy the newly developed electromagnetic Darwin version of the two-dimensional particle code EDIPIC-2D. The electrostatic version of EDIPIC-2D code has already been used for several applications and is benchmarked and validated [76, 29, 30, 77, 78, 79, 80]. In order to simulate an inductively coupled discharge, we implemented the electromagnetic Darwin scheme as detailed in Ref. [81]. For ease of analysis, the simulations are performed in Cartesian geometry. Figure 1 (a2) and (b2) depict the simulation domain. The rectangular chamber dimensions are $D_{x}\times D_{y}=80~{}\mathrm{mm}\times 80~{}\mathrm{mm}$, with a dielectric slab located at $70~{}\mathrm{mm}<y<80~{}\mathrm{mm}$ with antennas positioned within the dielectric (indicated by black rectangles showing them in cross-section) and the plasma occupying the rest of the simulation domain. Because the code implements a direct-implicit time advance [82, 71, 9, 81], we set the cell size to be $\Delta x=0.33~{}\mathrm{mm}$, resulting in overall grid dimensions of $N_{x}\times N_{y}=240\times 240$. The antenna currents are $180^{\circ}$ out of phase, oscillating with the frequency $f=\omega/2\pi$ in the range of $1$ to $10~{}\mathrm{MHz}$ and the amplitude $I_{coil}=60~{}\mathrm{to}~{}280~{}\mathrm{A}$. The amplitude of the coil current is relatively high because it is necessary to sustain the the plasma. Similar current values have been used in experimental ICP systems [25, 26, 83]. The domain is surrounded by conductive boundaries and secondary emission is neglected. Initially, the system contains Argon gas at the pressure $p=5~{}\mathrm{mTorr}$ and a plasma with a uniform density $N_{e}=N_{i}=10^{17~{}}\mathrm{m}^{-3}$, electron temperature $T_{e}=2.0~{}\mathrm{eV}$, and ion temperature $T_{i}=0.03~{}\mathrm{eV}$. $1000$ macro-particles per cell are specified initially for all species. Each simulation is analyzed after reaching a steady state.
Figure 1: Time evolution of physical quantities and steady-state profiles. Panel (a) is for the low frequency case $f=1\mathrm{MHz}$ and panel (b) is for the $f=10~{}\mathrm{MHz}$ case. Subplots (a1) and (b1) show time evolution of the ion density and electron temperature. The probes are placed at the center, shown by black stars in (a3) and (b3). Subplots (a2) and (b2) show 2D maps of the magnetic field strength with superimposed magenta curves tracing the field lines. Black rectangles are the cross-sections of the coil wires, and the black horizontal line is the plasma facing boundary of the dielectric. The profiles in (a2) and (b2) are plotted at the phase when the RF magnetic field is at the maximum. The profiles in (a3)-(b4) are the time averaged ion density and energy deposition, where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes time average.
Results–We first focus on two cases, with $I_{coil}$ at $130A$ and the frequencies of $f=1~{}\mathrm{MHz}$ and $f=10~{}\mathrm{MHz}$. Figure 1 shows the time histories and spatial profiles of the plasma for these cases. The time evolution of the ion density and electron temperature shown in Fig. 1 (a1) and (b1) indicates that the simulations have reached a steady state. Subplots (a2) and (b2) show the magnetic field lines with superimposed color maps of the field strength. We see that the magnetic field amplitude is almost the same between the low- and the high-frequency cases, because it depends mainly on the coil current ($B\propto I_{coil}$). Comparing the ion density profiles between Fig. 1 (a3) and (b3), we see a difference by nearly two orders of magnitude . This is due to a comparable difference in energy deposition between (a4) and (b4), because $E_{z}\cdot J_{ze}\propto\omega^{2}I_{coil}$.
A close examination of the current profiles for the two cases shown in Fig. 2 (a1) and (b1) reveals two notable distinctions. First, a pronounced jet-like structure of the electron current is observed in the time history for the low frequency case. The jet-like structure propagates into the plasma at around $t=29.5\mu s$, with a speed close to the electron thermal velocity $v_{th}$.
The second distinction is in the sample particle trajectories shown in Fig. 2 (a2) and (b2) for low and high frequency cases, respectively. For low frequency case, the electrons are magnetized with a gyro-radius ($\sim 2\mathrm{mm}$) smaller than the skin depth ($\sim 10~{}\mathrm{mm}$) and follow the RF magnetic field line. The electrons are, however, not magnetized in the high frequency case. To track the electrons, we randomly sampled 400 particle trajectories originating within the cell at $x=68.0~{}\mathrm{mm}$ and $y=65.7~{}\mathrm{mm}$ (the location marked by the black stars in Fig. 1 (a4) and (b4)) and at the time step near $t=29.2~{}\mu s$. The reason for choosing this location is because in the $1~{}\mathrm{MHz}$ case, most new electrons are produced by ionization near the coil (see Fig. 3 in our accompanied paper [84]). Among all the particles that were tracked in the simulation of the $f=1~{}\mathrm{MHz}$ case, around $80\%$ followed trajectories similar to the one plotted in Fig. 2 (a2). A movie showing the time evolution of physical quantities at the electron’s location is available in the supplemental material [85].
Figure 2: Subplots (a1) and (b1) show the color maps of the current density vs. $y$ and $t$; the domain cross sections are taken along the black dashed lines in Fig. 1 (a4) and (b4), respectively. Subplots (a2) and (b2) show representative particle trajectories, respectively, for the $1~{}\mathrm{MHz}$ and $10~{}\mathrm{MHz}$ cases over the time intervals marked in subplots (a3) and (b3). The red arrows in (a2) designate the counteracting electrostatic and Lorentz forces acting upon the test particle in the $1~{}\mathrm{MHz}$ case. Subplots (a3) and (b3) show the time evolution of the particle velocity and the drift velocity components, with $v_{\nabla B}$ being the drift due to the magnetic field gradient, $v_{curvB}$ being the curvature drift, and $v_{E_{y}\times B}$ being the $E_{y}\times B$ drift. The vector potential, expressed as $eA/m_{e}$, and the canonical momentum, $v_{z}+eA/m_{e}$ for the $1~{}\mathrm{MHz}$ case are traced by the green and light blue lines, respectively.
The above behavior is in significant contrast to the findings of Ref. [70], where the Larmor radius was assumed to be much larger than the skin depth. It is evident from Fig. 2 (a2) that electrons are magnetized and follow the magnetic field line. The oscillations of $v_{y}$ and $v_{x}$ shown in (a3) indicate that the electron undergoes a cyclotron motion in $y$ direction and bounces parallel to the magnetic field in $x$ direction. It is important to note that the electrons’ $y$ locations remain almost unchanged. This is due to the plasma generated ambipolar electric field, $E_{y}$, caused by the charge separation driven by the magnetic part of the Lorentz force $\vec{v}_{z}\times\vec{B}$ acting on the electrons [86]; the $E_{y}$ field counterbalances the magnetic part of the Lorentz force, to make the time-averaged net force on the electrons in $y$ direction vanish. Figure 2 (a3) further shows that the particle motion in $z$ direction is due to particle drifts, with $E_{y}\times B$ drift being the dominant one. As the RF magnetic field is $\pi/2$ out of phase with the inductive electric field and it decreases as the electric field increases, the electrons eventually demagnetize near the phase when the magnetic field becomes small: the electrons stop following the magnetic field lines and move into the plasma interior at around $t=29.5~{}\mu\text{s}$, forming a jet-like structure shown in Fig. 2 (a1). Note that during this process, the electron’s canonical momentum $m_{e}v_{z}+eA$ is conserved (denoted by the light blue line in Fig. 2 (a3)), where $A$ is the vector potential. A sample test particle in the $f=10~{}\mathrm{MHz}$ case, however, travels over the entire simulation domain in the $10$ RF periods shown in the figure, interacting with the coil field in a transient way, as described in multiple previous works; see, e.g., Refs. [68, 69].
To better understand the representative electron trajectory shown in Fig. 2 (a3), we turn to a simplified 1D model of electron motion along the vertical dashed line in Fig. 1 (a4)). Note that the electron motion parallel to magnetic field is not considered because the bounce parallel to the magnetic field occurs with lower frequency, and is not expected to significantly affect the electron current. Here, the simulation data shows that the vector potential can be approximated as $A_{z}(y,t)=A_{max}\exp(-y/\delta)cos(\omega t)$ (where $\delta$ is the skin depth). The RF magnetic field in the $x$ direction is $B_{x}(y,t)=\partial A_{z}/\partial y$ and the RF inductive electric field in the $z$ direction is $E_{z}(y,t)=-\partial A_{z}/\partial t$. The electron equation of motion reads
$m_{e}\frac{dv_{y}}{dt}=qv_{z}B_{x}-q\frac{\partial\Phi}{\partial y},\quad m_{e}\frac{dv_{z}}{dt}=qE_{z}-qv_{y}B_{x},$ (1)
where $\Phi$ is the in-plane electrostatic potential of the ambipolar electric field $E_{y}(y,t)$ generated by plasma. In this model, the electrons move in the $y-z$ plane under the influence of both the electrostatic and inductive fields. They do not undergo pure cyclotron motion with only the RF magnetic field $B_{x}(y,t)$. Since the canonical momentum $m_{e}v_{z}+qA_{z}=const$ is seen to be conserved, the above equations reduce to [68, 69]
$\frac{d^{2}y}{dt^{2}}=\frac{q}{m_{e}}\frac{\partial A_{z}}{\partial y}\left(v_{z0}-\frac{q}{m_{e}}(A_{z}-A_{z0})\right)-\frac{q}{m_{e}}\frac{\partial\Phi}{\partial y}\equiv-\frac{q}{m_{e}}\frac{\partial U}{\partial y},$ (2)
where $A_{z0}$ is the vector potential at the initial time and position of the electron being considered and the effective potential, $U(y,t)$, has been introduced as
$U=q\frac{A^{2}_{z}}{2}-(qA_{z0}+m_{e}v_{z0})A_{z}+\Phi.$ (3)
[…]
Theory– Although previous theory [68, 69] described many aspects of the low frequency ICP regime, there are new effects not addressed in these papers, notably the formation of jet-like currents and the resulting power deposition (see our accompanied paper Ref  [84] for specifics). These findings are connected to the fact that, in a realistic ICP discharge and indeed as we observe in simulations, a large amount of electrons is produced by ionization near the coil and, as shown above, their motion can be treated as a cyclotron rotation around a guiding center combined with a drift in $z$ direction. To obtain an analytical expression for the electron current, we find the velocity distribution as a solution to the Vlasov equation obtained by integration along unperturbed particle trajectories. Two major assumptions are involved: 1) The maximum electron bounce frequency $\Omega_{max}$ is much higher than the driving frequency, $\omega$. Indeed, for $1~{}\mathrm{MHz}$ case $\Omega_{max}\sim 10^{8}~{}\mathrm{s}^{-1}$ and $\omega=6.28\times 10^{6}~{}\mathrm{s}^{-1}$. 2) The plasma density and temperature gradients in the $y$ direction are much larger than those in the $x$ direction and we only assume a 1D spatial variation, specifically along the black dashed line in Fig. 1 (a4). Similar assumptions have been made in the pioneering paper of Tuszewski [17]. Note that the theory outlined below only works in the new physical regime and we use it to obtain an approximate analytical solution for the electron current. Under these stated assumptions and caveats, we first write the Vlasov equation for electrons as
$\frac{D}{Dt}f=\frac{\partial f}{\partial t}+\vec{v}\cdot\frac{\partial f}{\partial\vec{r}}+\frac{q}{m_{e}}(\vec{v}\times\vec{B}+\vec{E})\cdot\frac{\partial f}{\partial\vec{v}}=0.$ (4)
To arrive at an analytical solution for the current, the distribution function $f$ is split into a background Maxwellian $f_{0}$ with local electron density and temperature as its parameters; and $f_{1}$, a perturbation to it. Therefore, the kinetic equation for the perturbed electron distribution function $f_{1}$ is
$\frac{\partial f_{1}}{\partial t}+\vec{v}\cdot\frac{\partial f_{1}}{\partial\vec{r}}+\frac{q}{m}(\vec{v}\times\vec{B}+\vec{E}_{sc})\cdot\frac{\partial f_{1}}{\partial\vec{v}}=\frac{q}{m}\frac{\partial\vec{A}}{\partial t}\cdot\frac{\partial f_{0}}{\partial\vec{v}},$ (5)
where $\vec{A}$ is the vector potential of the EM field and $E_{sc}$ is the longitudinal electric field $E_{y}$ resulting from charge separation. Note that we have $\vec{E}_{sc}\gg\partial\vec{A}/\partial t$, based on simulations. Essentially, we consider the case where the RF magnetic field strongly alters the particle trajectory. This is the opposite case to kinetic theories describing the high frequency limit, see e.g., Ref. [45], where the RF magnetic field is neglected. Electron-neutral collisions are neglected because the collision frequency $\nu_{en}=3.2\times 10^{6}~{}\mathrm{s}^{-1}<\omega\ll\Omega_{max}$. Accounting for collisions is straightforward and we have verified that these weak collisions do not affect our results in a significant manner. The full derivation is presented in our accompanied paper [84]. The total current in $z$ direction is [87, 88]
[…]
where $\vec{v}_{z,d}$ is the drift velocity in $z$ direction and $\vec{v}_{\perp}$ is the perpendicular gyro-velocity. Upon integrating Eq. (5) along the unperturbed electron trajectory and making some justified assumptions, we obtain an analytical expression for $\hat{f}_{1}$ (the Fourier transform of $f_{1}$). The approximation to the electron current $J_{ze,1}$ obtained by integrating $\hat{f}_{1}$ is then
$J_{ze,1}(t)=Re\left\{\sum_{k_{y}}\hat{J}_{ze,1}(t)e^{ik_{y}y}\right\}=Re\left\{\sum_{k_{y}}\left[\frac{\omega q^{2}N_{e}\hat{A}(t)}{T_{e}}v^{\prime}_{d}\frac{v_{F}}{\omega}e^{-\xi}I_{0}(\xi)-\right.\right.\\ \left.\left.i\frac{\omega q^{2}N_{e}\hat{A}(t)}{T_{e}}\frac{(v^{\prime}_{d}+v_{F})k_{y}v^{2}_{th}}{\Omega\omega}e^{-\xi}\left(I_{0}(\xi)-I^{\prime}_{0}(\xi)\right)-\right.\right.\left.\left.\frac{\omega q^{2}N_{e}\hat{A}(t)}{T_{e}}\frac{v^{2}_{th}}{\omega}e^{-\xi}\left(2\xi I_{0}(\xi)-2\xi I^{\prime}_{0}(\xi)\right)\right]e^{ik_{y}y}\right\},$ (7)
where $Re\left\{...\right\}$ denotes the real part, $\hat{J}_{ze,1}$ and $\hat{A}$ denote the Fourier components, $N_{e}$ is the time-averaged local electron density, $T_{e}$ is the time-averaged local electron temperature, both from definition of $f_{0}$. Also, $\Omega(t)\approx qB(t)/m_{e}$, $v_{F}\approx-\langle F\rangle/qB(t)+E_{y}/B(t)$ is the total electron drift velocity (where $E\times B$ drift is dominant), ${\langle F\rangle}=-\langle m_{e}(v^{2}_{\perp}/2+v^{2}_{||})\rangle\nabla_{\perp}lnB$ is the average force (over particles) due to the magnetic field gradient and curvature, $v_{\perp}$ and $v_{||}$ are the gyro-velocity components in cylindrical coordinates, $I_{0}$ is the $0\mathrm{{}^{\prime}s}$ order modified Bessel function of the first kind (the prime on it denotes derivative), $\xi=(k_{y}v_{th}/\Omega(t))^{2}$, where $k_{y}$ is the argument of the Fourier component of ${A}$, and the diamagnetic drift velocity is $v^{\prime}_{d}\approx T_{e}/qB(t)dlnN_{e}/dy$, accounting for the density gradient. The first term in Eq. (7) is due to the drift $\vec{v}_{z,d}$ whereas the second and third terms represent contribution from $\vec{v}_{\perp}$, which is the response to the RF electromagnetic field modified by gyro-motion. The guiding center current comes mainly from the $E_{y}\times B$ drift in $f_{0}$ as
$J_{ze,0}(t)=qN_{e}v^{\prime}_{F}\left(1-\exp\left(-\frac{1}{2}\frac{v^{2}_{\perp,tr}}{v^{2}_{th}}\right)\right),$ (8)
where $v_{\perp,tr}=qB(t)\delta/2m_{e}$ is the maximum perpendicular velocity for the electrons to be trapped in the effective potential well (see Eq. (9)), since only the trapped electrons in $f_{0}$ contribute to the current. The velocity $v^{\prime}_{F}=E_{y}/B(t)$ contains only the $E\times B$ drift. Note that in Eq. (8) we have neglected the diamagnetic current since in our case, it is much smaller than the $E\times B$ current [84].
Our case is more complicated than the standard gyrokinetic treatment [88, 89, 90] in two aspects. First, only the electrons are magnetized, so the current from electron $E_{y}\times B$ drift dominates and cannot be canceled with that of the ions. Second, the guiding magnetic field is time-varying and the drift velocity $v_{F}$ is comparable to or larger than the thermal velocity $v_{th}$, making the terms associated with $v_{F}$ dominant. Because of this, the contribution from $f_{1}$ to the current is not negligible because there are terms proportional to $N_{e}$ and $v_{F}$. In fact, $|J_{ze,1}|\sim 0.5|J_{ze,0}|$. Equation (7) for $J_{ze,1}$ contains terms involving Bessel functions. Those represent the Finite Larmor Radius (FLR) effect. Physically, the FLR effect occurs because when a particle undergoes gyromotion and a field gradient is present, the particle tends to spend more time on the low-field side than the high-field side [91, 92, 93, 89]. As a consequence, the effective field acting on the particle will be reduced, hence reducing the current. Therefore, in fact, $J_{ze,1}$ takes the sign opposite to that of $J_{ze,0}$ and acts to reduce the current. A reduced expression based on Eq. (6) is also provided in our accompanied paper (Eq. (12) of Ref. [84]). It can be used for estimating the amplitude of the current in fluid simulation codes and experimental work. Physically, the dominant influence of the RF magnetic field on particle motion causes the relationship between the electron current and the RF electric field to be nonlinear, indeed indicating the presence of a new physical regime. With the gradual increase of coil current, the plasma density increases and the electron trajectory changes to the classical type as seen in Fig. 2 (b3), with all drifts becoming negligible and the electron current taking the form predicted in existing kinetic theory $\hat{J}_{ze,old}=\omega q^{2}/m_{e}\int v_{z}/(\omega-k_{y}v_{y})\partial f_{0}/\partial\vec{v}\cdot\hat{A}d\vec{v}$ [16, 45, 84]. When neglecting all drift velocities and taking the limit $B\to 0$, one can also prove that the third term contributing to $J_{ze,1}$ in our new theory reduces to the previously known kinetic result [84].
Figure 4: Comparing PIC simulations with theory for (a): $1~{}\mathrm{MHz}$ case and (b): $10~{}\mathrm{MHz}$ case. Subplots (a1) and (b1) show the electron current $J_{ze}$ and subplots (a2) and (b2) show the energy deposition rate. The comparison is made at the locations denoted by the black stars in Fig. 1 (a4) and (b4). For our new kinetic theory, we use Eq. (6) with $k_{y}=2\pi n_{y}/L_{y}$ with $n_{y}$ in the range from $1$ to $210$ (the number of cells occupied by plasma) to compute $J_{ze,1}$. The fluid approach is from Ref. [17] and the non-local kinetic calculation is from Ref. [45].
Figure 4 compares the time evolution of $J_{ze}$ and of the energy deposition between the theoretical predictions and the PIC simulations. For our presently proposed theory (black) and the Tuszewski theory (red dashed, based on a fluid model without accounting for self-consistent electrostatic field) [17], we evaluate the electric field and other relevant quantities at the location marked by the black star in Fig. 1 (a4) and (b4). We then calculate the electron current from Eq. (6) and the conductivity given in Ref. [17], respectively. To carry out a calculation based on the previous non-local kinetic theory developed for high frequency ICP (blue), we additionally employ the code provided in Ref. [45] to perform a 2D simulation of the entire system, and then extract the electric field and electron current at the same location marked in Fig. 1 (a4), (b4). Note that this existing kinetic theory assumes plasma density to be uniform across the simulation domain. It is seen that the fluid theory does not predict the current and the energy deposition rate, whereas the previous kinetic theory performs well only in the $f=10~{}\mathrm{MHz}$ case. Our new theory, on the other hand, displays a much better agreement for both the current and the energy deposition at low frequency. We identify the new regime dominated by periodic bursty energy deposition as the “periodic burst regime”. The above comparison is performed at a single location because the electrons are trapped in the strong effective potential and the plasma response is nearly local. When averaging over the entire simulation domain, the jet-like current, comprised of the released trapped electrons (see Fig. 2 (a3)), contributes around $70\%$ to the total energy deposition.
Parameter Scan–Next, we perform a parameter scan over $I_{coil}$ and $\omega$ to identify the onset boundary for the periodic burst regime. Figure 5 shows the average energy deposition, average ion density and average electron temperature for different values of the driving waveform parameters. We see that at lower frequencies there is a distinct jump in energy deposition from $10^{-2}W/cm^{3}$ to $10^{0}W/cm^{3}$, indicating a transition from the periodic burst regime to the anomalous (non-local) skin effect regime. A similar jump is observed for the ion density. Physically, this transition occurs because in the periodic burst regime, the electric field and current are nearly out of phase during most of the RF period. This is because the strong RF magnetic field makes plasma response nearly local. Only in the phase of the electron jet the energy deposition is seen to become significant. Therefore, the effective energy deposition becomes much lower than in the anomalous skin effect regime, making the electron density also much lower. Therefore, based on this understanding of particle dynamics, the criterion for transition between the bursty regime and the usual anomalous non-local regime can be produced by equating twice the electron gyro-radius with the anomalous field skin depth:
$2\frac{m_{e}v_{th}}{qB}=\left(\frac{v_{th}c^{2}}{\omega\omega^{2}_{pe}}\right)^{1/3}.$ (9)
When the diameter of the electron gyro-circle is larger than the skin depth, the electrons are unlikely to undergo a full gyro-cycle, and are therefore no longer being trapped in the RF effective potential. Based on Eq. (9), an estimate of the plasma density at the threshold of the periodic burst regime is obtained. The cases with different driving frequencies will have different transition values for the plasma density, as indicted in Fig. 5 (b) by the dashed lines. One expects to obtain the periodic burst regime in the region located below the dashed lines. A reasonable agreement is found, with the red and blue stars in (b) denoting the transition point, indicating the validity of Eq. (9) for finding the boundary between the periodic burst regime and the anomalous non-local regime.
Figure 5: The solid lines show, in a steady state, (a) energy deposition rates, averaged over the entire simulation domain and RF period, (b) time-averaged ion density at the probe location denoted by the black star in Fig. 1 (a4), and (c) time-averaged electron temperature at the probe location denoted by the black star in Fig. 1 (a3), for different values of the coil current and driving frequency. The dashed lines denote the onset threshold for the periodic burst regime, which follows the scaling law of $N_{e}\sim I_{coil}^{3}/\omega T_{e}$ according to Eq. (9). Each dashed line denotes the boundary for same frequency as the solid line of the same color in (b). As an estimate, the relevant magnetic field and temperature data for calculating the density (denoted by dashed lines) based on Eq. (9) are evaluated at the black star locations in Fig. 1 (a4) and (b4). The red and blue stars in (b) denote the transition points between the two regimes.
Conclusions–We identify a periodic burst energy deposition regime in low-frequency ICP discharges where the electrons are magnetized and their dynamics is dominated by bouncing in the effective potential of the electromagnetic and electrostatic fields through most of the RF period, so that the electron response becomes nearly local in space. When the RF magnetic field becomes weak, the electrons demagnetize, forming a jet-like current which propagates quickly into the plasma and accounts for a sizable fraction of the energy deposition rate. A new kinetic theory is proposed to estimate the electron current, which is in good agreement with particle-in-cell simulations. The condition in the driving waveform parameter space for the transition from anomalous skin effect regime (unmagnetized electrons) to the periodic burst regime is identified as a result of analyzing the data produced in a series of 2D simulations and is also analytically derived. A significant jump in the plasma density is observed during such transition, akin to what was observed in the E-H mode transition [94, 95, 20, 83, 96]. Therefore, our finding also offers a new possible explanation to the density jumps in the transition between E-mode and H-mode in low-frequency, low pressure ICP. We propose an estimate for the amplitude of the current density in the skin layer (Eq. (6)) and a criterion for transition to the periodic burst regime (Eq. (9)) in a form that can be verified in future experiments.
Acknowledgement– The authors thank Dr. Edward Startsev for checking the derivations in this paper. The authors thank Dr. Sarvesh Sharma for the discussions of high frequency harmonics. The work was partially supported by the National Natural Science Foundation of China (Grant No. 12305223) and the National Natural Science Foundation of Guangdong Province (Grant No.2023A1515010762). This research was also supported in part by the U.S. Department of Energy, Office of Fusion Energy Science under Contract No. DE-AC02-09CH11466.