### v3

# Nonlinear particle motion and bursty periodic energy deposition in inductively coupled plasmas

Abstract
Two-dimensional, electromagnetic particle-in-cell simulations are employed to study particle kinetics and power deposition in inductively coupled plasmas. We show that in the regime of low frequency ($\sim\mathrm{MHz}$) and low pressure, the motion of electrons can be highly nonlinear in the skin region near the antenna coil. Through most of the RF cycle, the electrons are strongly magnetized, with energy deposition being small. However, for a brief period around the null of the oscillating magnetic field, the electrons demagnetize, causing a jet-like current penetrating into the bulk plasma. During these brief periods, the power deposition becomes high, exhibiting a periodic burst nature. Based on kinetic theory, we provide new analytical expressions for the plasma current and energy deposition. A criterion for transition between the newly identified low-frequency, periodic-burst regime and the usual anomalous non-local skin effect regime is proposed and verified using a series of fully kinetic 2D particle-in-cell simulations.
preprint: APS/123-QED
I
Introduction–Energy deposition into electrons by a time-varying electromagnetic field is the primary method of sustaining the plasma in gas discharges for material processing [1, 2], such as Capacitively Coupled Plasma (CCP) [3, 4, 5, 6, 7, 8, 6, 9, 10, 11] and Inductively Coupled Plasma (ICP) discharges [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Driven by industry demands to achieve atomic-scale precision, plasma processing systems tend to operate at low gas pressure (few mTorr) [24, 25, 26, 27, 28, 29, 7, 5, 30, 9, 31, 32, 33], making it crucial to understand the underlying electron dynamics [34, 35]. For low-pressure, high-frequency ($f=13.56~{}\mathrm{MHz}$ or higher) ICP discharges, the skin effect was found to be of the anomalous type [25, 26, 27, 28, 36, 29, 37, 38, 39, 40], meaning that the skin depth, $\delta$, is determined by non-local kinetics as $\delta\sim(v_{th}c^{2}/\omega\omega^{2}_{pe})^{1/3}$ instead of $\delta\sim c/\omega_{pe}$ for the normal skin effect, where $v_{th}$ is electron thermal velocity, $c$ is the speed of light, $\omega$ is the driving frequency, and $\omega_{pe}$ is the electron plasma frequency. Electron heating in the anomalous regime is mainly due to wave-particle resonance of low-energy electrons, which can be treated similarly to the Landau damping [16, 41, 42, 40, 43, 18, 34, 44]. The anomalous regime was also called the non-local regime because over one RF period the electrons travel a significant distance $d\sim v_{th}/\sqrt{\omega^{2}+\nu^{2}}>\delta$ [40], where $\nu$ is the collision frequency. As a result, the plasma current is a non-local function of the RF electric field. Multiple modeling studies have been performed for the ICP discharges, utilizing the non-local kinetic approach [16, 45, 34, 46, 47, 48, 49, 50] to predict the spatial distribution of the plasma current and power deposition.
Although historically ICP discharge systems operated mostly at high frequency [51, 52], it is recently shown that operating at lower frequency ($\sim 1$MHz) presents certain advantages [53, 54, 55, 56]. At low frequency, the capacitive coupling with the plasma can be reduced [57], the external circuit losses decrease [58, 57], and higher power transfer efficiency can be achieved [54], thereby improving the performance of plasma processing systems. More complicated ICP operations under pulsed waveforms [59, 60, 61, 62, 63, 64, 65] also require a fundamental understanding of ICP discharge at low frequency. Low-frequency ICP discharges are also utilized as sources of negative ions for generating fast atomic beams in fusion applications [66, 67, 68, 69].
Due to the lack of self-consistent kinetic 2D modeling, the physics of low-frequency ICP discharges is not sufficiently understood, because in experiments it is difficult to perform detailed measurements of plasma properties and electromagnetic fields within the skin layer and correspondingly deduce particle kinetics [55, 36]. Although it was previously shown that at low frequencies the particle dynamics could become highly nonlinear [70, 71], those studies were not based on fully self-consistent kinetic simulations and used simplified test-particle analysis. Ref. [72] classified ICP discharges according to their estimated skin depth, assuming that the electron gyro-radius is much larger than the skin depth. However, as will be shown, it is not the case at high coil current where the RF magnetic field strongly affects particle motion. Existing kinetic simulation studies have mostly been limited to one-dimensional [73, 74, 75, 46] or high-frequency ICP discharges [76, 77]. More importantly, to our knowledge, a theory that could explain the particle dynamics and predict the analytical expression for the electron current in a low-frequency inductive discharge has been lacking. Therefore, a comprehensive numerical and analytical study is necessary to understand the relevant particle dynamics and the resulting power deposition in low-frequency ICP discharges.
We demonstrate that a strongly nonlinear skin effect regime with periodic bursty energy deposition presents at low RF frequency, where electrons are magnetized and power deposition is low during most of the RF period. Demagnetization occurs when the RF magnetic field passes through zero, and electrons quickly escape the area near antennas to form a current jet, causing periodic bursts in the energy deposition. A new kinetic theory is developed and predicts well the plasma current in this bursty regime. To our knowledge, this work presents the first comprehensive 2D Particle-In-Cell (PIC) simulations of a low frequency ICP discharge.
Method – To simulate inductively coupled discharge, we implemented the electromagnetic Darwin scheme to EDIPIC-2D [78, 79, 31, 32, 80, 81, 82, 83]. The simulations are performed in Cartesian geometry. Figure 1 (a2) and (b2) depict the simulation domain. The rectangular chamber dimensions are $D_{x}\times D_{y}=80~{}\mathrm{mm}\times 80~{}\mathrm{mm}$, with a dielectric slab located at $70~{}\mathrm{mm}<y<80~{}\mathrm{mm}$ and antennas positioned within it (black rectangles showing cross-section). The plasma occupies the rest of the simulation domain. Because the code implements a direct-implicit time advance [84, 73, 9, 78], we set the cell size to be $\Delta x=0.33~{}\mathrm{mm}$, resulting in overall grid dimensions of $N_{x}\times N_{y}=240\times 240$. The antenna currents are $180^{\circ}$ out of phase, oscillating with the frequency $f=\omega/2\pi$ in the range of $1$ to $10~{}\mathrm{MHz}$ and the amplitude $I_{coil}=60~{}\mathrm{to}~{}280~{}\mathrm{A}$. The amplitude of the coil current is relatively high because it is necessary to sustain the plasma. Similar values have been used in experiments [27, 28, 85]. The domain is surrounded by conductive boundaries with secondary emission neglected. Initially, the system contains Argon gas at the pressure $p=5~{}\mathrm{mTorr}$ and a plasma with a uniform density $N_{e}=N_{i}=10^{17~{}}\mathrm{m}^{-3}$, electron temperature $T_{e}=2.0~{}\mathrm{eV}$, and ion temperature $T_{i}=0.03~{}\mathrm{eV}$. $1000$ macro-particles per cell are specified initially for all species. Each simulation is analyzed after reaching a steady state.
Figure 1: Time evolution of physical quantities and steady-state profiles. Panel (a) is for the low frequency case $f=1\mathrm{MHz}$ and panel (b) is for $f=10~{}\mathrm{MHz}$ case. Subplots (a1) and (b1) show time evolution of the ion density and electron temperature. The probes are placed at the center, shown by black stars in (a3) and (b3). Subplots (a2) and (b2) show 2D maps of the magnetic field strength with superimposed magenta curves tracing the field lines. Black rectangles are the cross-sections of the coil wires, and the black horizontal line is the plasma facing boundary of the dielectric. Subplots (a2) and (b2) are plotted at the phase when the RF magnetic field is at the maximum. Subplots (a3)-(b4) are the time-averaged (denoted by delimited-⟨$\langle$..$\rangle$ … ⟩) ion density and energy deposition.
Results–We first focus on two cases, with $I_{coil}$ at $130A$ and the frequencies of $f=1~{}\mathrm{MHz}$ and $f=10~{}\mathrm{MHz}$. Figure 1 shows the time histories and spatial profiles of the plasma for these cases. The time evolution of the ion density and electron temperature shown in Fig. 1 (a1) and (b1) indicates that the simulations have reached a steady state. Subplots (a2) and (b2) show the magnetic field lines with superimposed color-maps of the field strength. The magnetic field amplitude is almost the same between the low- and the high-frequency cases because it depends mainly on the coil current ($B\propto I_{coil}$). Comparing the ion density profiles between Fig. 1 (a3) and (b3), we see a significant difference. This is due to a comparable difference in energy deposition between (a4) and (b4), because $E_{z}\cdot J_{ze}\propto\omega^{2}I_{coil}$.
A close examination of the two cases shown in Fig. 2 reveals two notable distinctions. First, a pronounced jet-like structure of the electron current is observed in the time history of the low-frequency case (see Fig. 2 (a1)). The jet-like structure propagates into the plasma at around $t=29.5\mu s$, with a speed close to the electron thermal velocity $v_{th}$. The second distinction is in the representative particle trajectories shown in Fig. 2 (a2) and (b2) for low- and high-frequency cases. For low-frequency case, electrons are magnetized with a gyro-radius ($\sim 2\mathrm{mm}$) smaller than the skin depth ($\sim 10~{}\mathrm{mm}$). The electrons are, however, not magnetized in the high-frequency case. To track the electrons, we randomly sampled 400 particle trajectories originating from $x=68.0~{}\mathrm{mm}$ and $y=65.7~{}\mathrm{mm}$ (the location marked by the black stars in Fig. 1 (a4) and (b4)) at $t=29.2~{}\mu s$. We choose this location and time because in the $1~{}\mathrm{MHz}$ case, most new electrons are produced by ionization near the coil at this phase of the RF period (see Fig. 3 in our accompanied paper [86]). A movie for this particle is available in the supplemental material [87].
Figure 2: Subplots (a1) and (b1) show the color maps of the current density vs $y$ and $t$; the domain cross sections are taken along the black dashed lines in Fig. 1 (a4) and (b4), respectively. Subplots (a2) and (b2) show representative particle trajectories, respectively, for the $1~{}\mathrm{MHz}$ and $10~{}\mathrm{MHz}$ cases over the time intervals marked in subplots (a3) and (b3). The red arrows in (a2) designate the counteracting electrostatic and Lorentz forces acting upon the particle in the $1~{}\mathrm{MHz}$ case. Subplots (a3) and (b3) show the time evolution of the particle velocity and the drift velocity components, with $v_{\nabla B}$ being the drift due to the magnetic field gradient, $v_{curvB}$ being curvature drift, and $v_{E_{y}\times B}$ being $E_{y}\times B$ drift. The vector potential, expressed as $qA/m_{e}$, and the canonical momentum, $v_{z}+qA/m_{e}$ for $1~{}\mathrm{MHz}$ case are traced by the green and light blue lines.
This particle behavior is in significant contrast to the findings of Ref. [72], where the Larmor radius was assumed to be much larger than the skin depth. It is evident from Fig. 2 (a2) that electrons are magnetized and follow the magnetic field line. The oscillations of velocities shown in (a3) indicate that the electron undergoes a cyclotron motion in $y-z$ plane and bounces parallel to the magnetic field in $x$ direction. It is important to note that the electrons’ $y$ locations remain almost unchanged. This is due to the plasma-generated ambipolar electric field [88], $E_{y}$, counterbalancing the Lorentz force $\vec{v}_{z}\times\vec{B}$ and making the time-averaged net force on the electrons in $y$ direction vanish. Figure 2 (a3) further shows that the particle motion in $z$ direction is due to particle drifts, with $E_{y}\times B$ drift being dominant [86]. As the RF magnetic field is $\pi/2$ out of phase with the inductive electric field and it decreases as the electric field increases, the electrons eventually demagnetize near the phase when the magnetic field becomes small: the electrons move into the plasma interior at around $t=29.5~{}\mu\text{s}$, forming a jet-like structure shown in Fig. 2 (a1). During this process, the electron’s canonical momentum $m_{e}v_{z}+qA$ is conserved (denoted by the light blue line in Fig. 2 (a3)), where $A$ is the vector potential. A sample test particle in the $f=10~{}\mathrm{MHz}$ case, however, travels over the entire simulation domain, interacting with the coil field in a transient way, as described in multiple previous works [70, 71].
Theory–As shown in Fig. 2 (a3) and justified in Appendix A and B of Ref. [86], electron motion can be treated as a cyclotron rotation in $y-z$ plane around a guiding center drifting in $z$ direction. To obtain an analytical expression for electron current, the electron distribution function $f$ is split into a background Maxwellian $f_{0}$ as function of local electron density and temperature; and $f_{1}$, a perturbation to it. Two major assumptions are involved: 1) The maximum electron cyclotron frequency $\Omega_{max}$ is much higher than the driving frequency, $\omega$. Indeed, for $1~{}\mathrm{MHz}$ case $\Omega_{max}\sim 10^{8}~{}\mathrm{s}^{-1}$ and $\omega=6.28\times 10^{6}~{}\mathrm{s}^{-1}$. 2) The plasma density gradient in the $y$ direction is much larger than that in $x$ direction and we only consider a 1D spatial variation, specifically along the black dashed line in Fig. 1 (a4). Similar assumptions have been made in the pioneering paper of Tuszewski [17]. The theory below only works in the new physical regime and we use it to obtain approximate analytical solution for the electron current. The kinetic equation for the perturbed electron distribution function $f_{1}$ is
$\frac{\partial f_{1}}{\partial t}+\vec{v}\cdot\frac{\partial f_{1}}{\partial\vec{r}}+\frac{q}{m}(\vec{v}\times\vec{B}+\vec{E}_{sc})\cdot\frac{\partial f_{1}}{\partial\vec{v}}=\frac{q}{m}\frac{\partial\vec{A}}{\partial t}\cdot\frac{\partial f_{0}}{\partial\vec{v}},$ (1)
where $\vec{A}$ is the vector potential of the EM field and $E_{sc}$ is the longitudinal electric field $E_{y}$ resulting from charge separation. Note that we have $\vec{E}_{sc}\gg\partial\vec{A}/\partial t$ based on simulations. Essentially, we consider the case where the RF magnetic field strongly alters the particle trajectory. This is the opposite case to kinetic theories describing the high frequency case [47]. Electron-neutral collision is neglected because the collision is weak $\nu_{en}=3.2\times 10^{6}~{}\mathrm{s}^{-1}<\omega\ll\Omega_{max}$ and does not affect our results significantly. The full derivation is presented in our accompanied paper [86]. The total current in $z$ direction is [89, 90]
[…]
where $\vec{v}_{z,d}$ is the drift velocity in $z$ direction and $\vec{v}_{\perp}$ is the perpendicular gyro-velocity. Upon integrating Eq. (1) along the unperturbed electron trajectory, we obtain an analytical expression for $\hat{f}_{1}$ (the Fourier transform of $f_{1}$). The approximation to the electron current $J_{ze,1}$ obtained by integrating $\hat{f}_{1}$ is
$J_{ze,1}(t)=Re\left\{\sum_{k_{y}}\hat{J}_{ze,1}(t)e^{ik_{y}y}\right\}=Re\left\{\sum_{k_{y}}\left[\frac{\omega q^{2}N_{e}\hat{A}(t)}{T_{e}}v^{\prime}_{d}\frac{v_{F}}{\omega}e^{-\xi}I_{0}(\xi)-\right.\right.\\ \left.\left.i\frac{\omega q^{2}N_{e}\hat{A}(t)}{T_{e}}\frac{(v^{\prime}_{d}+v_{F})k_{y}v^{2}_{th}}{\Omega\omega}e^{-\xi}\left(I_{0}(\xi)-I^{\prime}_{0}(\xi)\right)-\right.\right.\left.\left.\frac{\omega q^{2}N_{e}\hat{A}(t)}{T_{e}}\frac{v^{2}_{th}}{\omega}e^{-\xi}\left(2\xi I_{0}(\xi)-2\xi I^{\prime}_{0}(\xi)\right)\right]e^{ik_{y}y}\right\},$ (3)
where $Re\left\{...\right\}$ denotes the real part, $\hat{J}_{ze,1}$ and $\hat{A}$ denote the Fourier components, $N_{e}$ and $T_{e}$ are the time-averaged local electron density and temperature, both from definition of $f_{0}$. Also, $\Omega(t)\approx qB(t)/m_{e}$, $v_{F}\approx-\langle F\rangle/qB(t)+E_{y}/B(t)$ is the total electron drift velocity (where $E\times B$ drift is dominant), ${\langle F\rangle}=-\langle m_{e}(v^{2}_{\perp}/2+v^{2}_{||})\rangle\nabla_{\perp}lnB$ is the average force (over particles) due to the magnetic field gradient and curvature, $v_{\perp}$ and $v_{||}$ are the gyro-velocity components in cylindrical coordinates, $I_{0}$ is the $0\mathrm{{}^{\prime}s}$ order modified Bessel function of the first kind (the prime on it denotes derivative), $\xi=(k_{y}v_{th}/\Omega(t))^{2}$, where $k_{y}$ is the argument of the Fourier component of ${A}$, and the diamagnetic drift velocity is $v^{\prime}_{d}\approx T_{e}/qB(t)dlnN_{e}/dy$. The first term in Eq. (3) is due to the drift $\vec{v}_{z,d}$, whereas the second and third terms represent contributions from $\vec{v}_{\perp}$. The guiding center current $J_{ze,0}$ related to $f_{0}$ comes mainly from $E_{y}\times B$ drift
$J_{ze,0}(t)=qN_{e}v^{\prime}_{F}\left(1-\exp\left(-\frac{1}{2}\frac{v^{2}_{\perp,tr}}{v^{2}_{th}}\right)\right),$ (4)
where $v_{\perp,tr}=|qB(t)|\delta/2m_{e}$ is the maximum perpendicular velocity for the electrons to be trapped by the magnetic field, since only the trapped electrons in $f_{0}$ contribute to the current. And $v^{\prime}_{F}=E_{y}/B(t)$. In Eq. (4) we have neglected the diamagnetic current since it is much smaller than the $E\times B$ current.
Our case is more complicated than the standard gyrokinetic treatment [90, 91, 92] in two aspects. First, only the electrons are magnetized, so the current from electron $E_{y}\times B$ drift dominates and cannot be canceled with that of the ions. Second, the guiding magnetic field is time-varying and the drift velocity $v_{F}$ is comparable to or larger than the thermal velocity $v_{th}$, making the terms associated with $v_{F}$ dominant. Because of this, the contribution from $f_{1}$ to the current is not negligible because there are terms proportional to $N_{e}$ and $v_{F}$. Equation (3) for $J_{ze,1}$ contains Bessel functions. Those represent the Finite Larmor Radius (FLR) effect. Physically, the FLR effect occurs because when a particle undergoes gyro-motion and a field gradient is present, the particle tends to spend more time on the low-field side than the high-field side [93, 94, 95, 91]. Therefore effectively, the field acting on the electrons will be reduced, hence reducing the current. Therefore, $J_{ze,1}$ takes the sign opposite to that of $J_{ze,0}$ and acts to reduce the current. The dominance of the RF magnetic field on particle motion significantly changes the electron current (see simplified formula Eq. (12) in Ref. [86]), indicating the presence of a new physical regime. With the gradual increase of coil current, the plasma density increases and the electron trajectory changes to the classical type as seen in Fig. 2 (b3), with all drifts becoming negligible and the electron current taking the form in existing kinetic theory $\hat{J}_{ze,old}=\omega q^{2}/m_{e}\int v_{z}/(\omega-k_{y}v_{y})\partial f_{0}/\partial\vec{v}\cdot\hat{A}d\vec{v}$ [16, 47].
Figure 3: Comparing PIC simulations with theory for (a): $1~{}\mathrm{MHz}$ case and (b): $10~{}\mathrm{MHz}$ case. Subplots (a1) and (b1) show the electron current $J_{ze}$ and subplots (a2) and (b2) show the energy deposition rate. The comparison is made at the locations denoted by black stars in Fig. 1 (a4) and (b4). For our new kinetic theory, we use Eq. (2) with $k_{y}=2\pi n_{y}/L_{y}$ with $n_{y}$ from $1$ to $210$ (number of cells occupied by plasma) to compute $J_{ze,1}$. The fluid approach is from Ref. [17] and the previous non-local kinetic calculation is from Ref. [47].
Figure 3 compares the time evolution of $J_{ze}$ and energy deposition between the theoretical predictions and the PIC simulations. For our new theory (black) and the Tuszewski theory (red dashed, based on fluid model) [17], we evaluate the electric field and other relevant quantities at the location marked by the black stars in Fig. 1 (a4) and (b4). We then calculate the electron current from Eq. (2) and the conductivity given in Ref. [17], respectively. To carry out a calculation based on the previous non-local kinetic theory developed for high frequency ICP (blue), we employ the code provided in Ref. [47] to perform a 2D simulation of the entire system, and then extract the electric field and electron current at the same location marked in Fig. 1 (a4), (b4). This existing kinetic theory assumes uniform plasma density. We see that the fluid theory does not predict the current and the energy deposition rate, whereas the previous kinetic theory performs well only for $f=10~{}\mathrm{MHz}$ case. Our new theory, on the other hand, displays a much better agreement for both the current and the energy deposition at low frequency. We identify the new regime dominated by periodic bursty energy deposition as the “periodic burst regime”. When averaging over the simulation domain, the jet-like current contributes around $70\%$ to the total energy deposition.
Parameter Scan–We perform a parameter scan over $I_{coil}$ and $\omega$ to identify the onset boundary for the periodic burst regime. Figure 4 shows the average energy deposition, ion density, and electron temperature for different values of the driving waveform parameters. We see that at lower frequencies there is a distinct jump in energy deposition, indicating a transition from the periodic burst regime to the anomalous (non-local) skin effect regime. A similar jump is observed for the ion density. This transition occurs because in the periodic burst regime, the electric field and current are nearly out of phase during most of the RF period, since the strong RF magnetic field makes plasma response nearly local. Only in the phase of electron jet will the energy deposition become significant. Therefore, the energy deposition becomes lower than in the anomalous skin effect regime, making the electron density also much lower. Based on this understanding of particle dynamics, the criterion for transition between the bursty regime and the anomalous non-local regime is estimated by equating twice the electron gyro-radius with the anomalous field skin depth [86]
$2\frac{m_{e}v_{th}}{|qB|}=\left(\frac{v_{th}c^{2}}{\omega\omega^{2}_{pe}}\right)^{1/3}.$ (5)
When the diameter of the electron gyro-circle is larger than the skin depth, the electrons are unlikely to undergo a full gyro-cycle, and are therefore no longer being confined by the magnetic field. Based on Eq. (5), an estimate of the plasma density at the threshold of the periodic burst regime is obtained. The cases with different driving frequencies will have different transition values for the plasma density, as indicated in Fig. 4 (b) by the dashed lines. One expects to obtain the periodic burst regime in the region located below the dashed lines. A reasonable agreement is found, indicating the validity of Eq. (5) for finding the boundary between the two regimes.
Figure 4: The solid lines show, in a steady state, (a) energy deposition rates, averaged over the entire simulation domain and RF period, (b) time-averaged ion density at the probe location denoted by the black star in Fig. 1 (a4), and (c) time-averaged electron temperature at the probe location denoted by the black star in Fig. 1 (a3), for different values of the coil current and driving frequency. The dashed lines denote the onset threshold for the periodic burst regime, which follows the scaling law of $N_{e}\sim I_{coil}^{3}/\omega T_{e}$ according to Eq. (5). Each dashed line denotes the boundary for the solid line of the same color in (b). The relevant magnetic field and temperature data for calculating the dashed lines is evaluated at the black stars in Fig. 1 (a4) and (b4). The red and blue stars in (b) denote the transition points between the two regimes.
Conclusions–We identify a periodic burst energy deposition regime in low-frequency ICP discharges where the electrons are magnetized, so that the electron response becomes nearly local during most of the RF period. When the RF magnetic field becomes weak, the electrons demagnetize, forming a jet-like current propagating quickly into the plasma and accounting for a sizable fraction of the energy deposition. A new kinetic theory is proposed to estimate the electron current, showing good agreement with PIC simulations. The condition for the transition from anomalous skin effect regime (unmagnetized electrons) to the periodic burst regime is both identified in simulations and given analytically. A significant jump in the plasma density is observed during such transition, akin to what was observed in the E-H mode transition [96, 97, 20, 85, 98]. Our finding also offers a new possible explanation for the density jumps in the transition between E-mode and H-mode in low-frequency, low pressure ICP. We propose an estimate for the current density in the skin layer (Eq. (2)) and a criterion for transition to the periodic burst regime (Eq. (5)) that can be verified in future experiments.
Acknowledgement– The authors thank Dr. Edward Startsev for checking the derivations in this paper. The authors thank Dr. Sarvesh Sharma for the discussions of high frequency harmonics. The work was partially supported by the National Natural Science Foundation of China (Grant No. 12305223) and the National Natural Science Foundation of Guangdong Province (Grant No.2023A1515010762). This research was also supported in part by the U.S. Department of Energy, Office of Fusion Energy Science under Contract No. DE-AC02-09CH11466.